OpenStudy (anonymous):

Interval notation 2 l3x+6l +12 =24

OpenStudy (anonymous):

isolate absolute values: I3x+6I=6

OpenStudy (anonymous):

then solve two different equations 3x+6=-6 and 3x+6=6 you should get two answers, if you need further assistance let me know

OpenStudy (anonymous):

x=-4 x= 4?